OpenStudy (anonymous):

I need help with an algebra problem. 1/5x+1/3y=19/5 15x+15y=15

OpenStudy (anonymous):

Here, you have a system of equations with two variables, and two equations. The easiest way to solve a problem like this with is to find a way to eliminate one variable. Lets try to eliminate x, since it is the first variable in the system. You can subtract one equation from the other, but this doesn't really help you unless one of the coefficient-variable combinations is the same in both equations. This would make one variable go away, allowing you to easily solve for the other variable. In order to make the coefficients of both x-variables equal, you have to multiply one equation by some constant to change the coefficient of x. In this case, you need to find a constant that, when multiplied by 1/5, equals 15. Then multiply the entire first equation by this constant, subtract one equation from the other, and you will easily be able to solve for y. Once you have solved for y, you can substitute this value back into either of the first equations, and solve for x as well. The final (x,y) solution will work for both equations.

OpenStudy (anonymous):

1/5x+1/3y=19/5 15x+15y=15 15x+15y=15------divide all by 15-----x+y=1---- which means X=1-Y u replace the above equation on the other 1 1/5x+1/3y=19/5---1/5(1-Y) +1/3Y=19/5-----Solve for Y & then After u get Y ----- X=1-Y --- So u get X do u get it

OpenStudy (anonymous):

Yes could you help me with another problem?

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

I have to determine if the line is parallel, perpendicular, or neither. 4x+y=-4 and 2y=10x-8

OpenStudy (anonymous):

SO for the 1st equation ---- 4x+y=-4 --- rearrange it --- Y =-4X-4 For the 2nd --- 2y=10x-8 ---rearrange divide everything by 2, it will be ---Y=5X-4 get 2 lines that pass through the equations for the 1st equation -- Y=-4X-4 -- When X=0---Y=-4 & another when Y=0---X=-1 So the lines that pass are --(X1,Y1)&(X2,Y2) (0,-4) & (-1,0) for the 2nd equation -- Y=5X-4---When X=0---Y=-4 & another when Y=0---X=4/5 So the lines that pass are --(X1,Y1)&(X2,Y2) (0,-4) & (4/5,0) So now find slope m1 =Y2-Y1/X2-X1-----0-(-4)/-1-0---=-4 m2 =Y2-Y1/X2-X1 ----0-(-4)/4/5-0)= 20/5 = 4 so its neither

OpenStudy (anonymous):

if the slope was 4 & -1/4 then perpendicular if the slope was 4 & 4then parallel

OpenStudy (anonymous):

@helpmewithmath2012 do u get it